Using the supermesh approach, find the current through each element of the

network below.


1
Expert's answer
2021-01-27T04:55:42-0500

"I 4\varOmega = i1 = -0.9 A"

"I 6\varOmega = i1- 3A = -3.9A"

"I 8\varOmega = i2- 3 = 4.1A"

"I 2\varOmega = i2 = 7.1A"